OpenStudy (anonymous):

four inces of heavy wet snow re equalvalent to an inch of rain. estimate thewater content in 10 inches of heavy wet snow

OpenStudy (saifoo.khan):

Heavy snow : Rain 4 : 1

OpenStudy (saifoo.khan):

Heavy snow : Rain 4 : 1 10 : x Cross multiply. 4(x) = 10(1) Solve for x.

OpenStudy (anonymous):

how do I cross multply

OpenStudy (anonymous):

would it be 4*10

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

maybe 4 inches

OpenStudy (anonymous):

\[\frac{ 4 }{ 1 } = \frac{ 10 }{ x }\] Cross multiply to solve for x

OpenStudy (anonymous):